You work for an ergonomic consulting company. Your group is in charge of assembling monitor arms for a new client. You have to assemble 73 of them in total. If your first one takes your group 41 minutes and you have a learning curve of 84%, how long do you expect the last unit will take?

A. 31.08 minutes

B. 13.93 minutes

C. 11.21 minutes

D. 9.78 minutes
Answer #1

We know,

Tn = T1*n^(ln(l)/ln(2)) [ Tn = Time taken to complete nth job,l = learning rate]

or, Tn = 41*73^(ln(0.84)/ln(2)) = 13.93432456 minutes

So correct answer is B. 13.93 minutes